OpenStudy (anonymous):

which choice is equivalent to the expression below? √- 20 A. -20i B. 2i √5 C. -2 √5 D. 2 √5i E. 20i

OpenStudy (ranga):

\[\sqrt{-20} = \sqrt{ -1 * 4 * 5} = \sqrt{-1} \times \sqrt{4} \times \sqrt{5}\]

OpenStudy (anonymous):

√-20 right ?

OpenStudy (ranga):

Simplify the right hand side. No. sqrt(-20) is what is given and you have to simplify. Look at the right hand side in my first reply. What is sqrt(-1)? What is sqrt(4)?

OpenStudy (anonymous):

√4 is 2 and the √ of -1 and 5 were decimals

OpenStudy (ranga):

What is square root of -1? sqrt(5) cannot be simplified and so leave it as it is.

OpenStudy (anonymous):

i thought u couldnt find the sqaure root of -1 cause its a negative

OpenStudy (ranga):

The symbol "i" is used to represent the square root of -1.

OpenStudy (anonymous):

i tried to find it for -1 but its says that its invalid

OpenStudy (ranga):

You cannot use a calculator to find sqrt(-1). sqrt(-1) = i where "i" is a complex or imaginary number and not a real number. i * i = i^2 = -1 i = sqrt(-1)

OpenStudy (anonymous):

im not sure on how to find it

OpenStudy (ranga):

You just use the symbol "i" in the place of sqrt(-1). You probably noticed "i" in some of your answer choices.

OpenStudy (anonymous):

yes i have but i still don't understand this

OpenStudy (ranga):

Look at the right hand side of my first reply. Replace sqrt(-1) with i Replace sqrt(4) with 2 Leave sqrt(5) as is. Multiply them all together. What do you get?

OpenStudy (anonymous):

-40 ?

OpenStudy (ranga):

\[\sqrt{-20} = \sqrt{ -1 * 4 * 5} = \sqrt{-1} \times \sqrt{4} \times \sqrt{5} = i * 2 * \sqrt{5}\]

OpenStudy (anonymous):

2i√5?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

so one is a imaginary number and the 5 and 4 came from the 20 ?

OpenStudy (ranga):

The letter "i" is an imaginary number not one. 1 on the right hand side of my first reply came form the negative sign which is equivalent to multiplying +20 and -1 which is -20.

OpenStudy (ranga):

For example, 12 = 4 * 3 -12 = -1 * 4 * 3

OpenStudy (anonymous):

then u would look for the √ for them

OpenStudy (ranga):

If you are asked to simplify the square root of a number, try to find the factors of the numbers. If any of the factors is a perfect square, then you can easily take the square root of that number. The number 4 is a perfect square whose square root is 2. 9, 16, 25 are all perfect squares. For example, Simplify sqrt(50) sqrt(50) = sqrt(25 * 2) = sqrt(25) * sqrt(2) = 5*sqrt(2) Simplify sqrt(-50) sqrt(-50) = sqrt(-1 * 25 * 2) = sqrt(-1) * sqrt(25) * sqrt(2) = i * 5*sqrt(2) = 5isqrt(2)

OpenStudy (anonymous):

wait, how do u knw wen a number is a real number ? like how do u tell ?

OpenStudy (ranga):

The numbers you are used to, positive numbers, negative numbers, zero, integers, rational numbers, irrational numbers are all real numbers. You can take the square root of any positive number and it will be a real number. But a square root of any negative number will be an imaginary number and square root of -1 has been assigned the symbol "i".

OpenStudy (ranga):

In the above example, square root of 50 is a real number: 5sqrt(2) square root of (-50) is an imaginary (or complex number): 5isqrt(2)